# Threat Modeling for AI: The Case for an Asset-Centric Approach

### v1

@@ I Introduction
This paper introduces an asset-centric methodology for threat modeling AI systems that addresses the unique security challenges posed by integrated AI agents. Unlike existing top-down frameworks that analyze individual attacks within specific product contexts, our bottom-up approach enables defenders to systematically identify how vulnerabilities—both conventional and AI-specific—impact critical AI assets across distributed infrastructures used to develop and deploy these agents. This methodology allows security teams to: (1) perform comprehensive analysis that communicates effectively across technical domains, (2) quantify security assumptions about third-party AI components without requiring visibility into their implementation, and (3) holistically identify AI-based vulnerabilities relevant to their specific product context. This approach is particularly relevant for securing agentic systems with complex autonomous capabilities. By focusing on assets rather than attacks, our approach scales with the rapidly evolving threat landscape while accommodating increasingly complex and distributed AI development pipelines.
@@ I-A AI-based Vulnerabilities
AI represents a shift in our computing paradigms towards data driven compute. Under traditional compute, a developer would create a program – i.e. logic rules – to generate a desired output from unseen input data. AI’s computing paradigm instead analyzes massive amounts of data to infer – i.e. learn or create – logic rules. These inferred rules are then used to generate a desired output when applied to unseen input data. The AI’s underlying decision making logic is, therefore, algorithmically inferred rather than explicitly encoded [9, 22]. This data driven nature is why AI’s decisions are often difficult to explain, and why AI is difficult to validate. Evasion, or jailbreak, attacks are an example of adversaries finding and exploiting corner cases or incorrectly correlated features within the AI’s inferred logic; i.e. the model. To compromise an AI’s decision, however, adversaries are not constrained to direct attacks on the model [28, 17, 23, 8]. Adversaries are also able to introduce vulnerabilities into an AI by affecting its training data, training process or algorithm, and validation process or goals. To augment their security policies against emerging AI-based vulnerabilities, defenders must reason about the many ways in which adversaries can affect their AI.
@@ I-C The case for a bottom-up perspective
Section II will introduce common types of AI Assets and describe their relationships. Section III will elaborate on the challenges defenders face when threat modeling AI. Section IV provides background on asset-centric threat modeling. Section V describes our bottom-up, asset-centric, approach for threat modeling AI. Sections VI and VII introduce Enterprise RAG and apply this analysis to it, as a case study. Section VIII describes existing frameworks for threat modeling AI and why they naturally complement an asset-centric approach. Section IX describes our ongoing work and next steps. Finally, Section X concludes.
II Relationships between AI Assets
The relationships between AI assets are inherent to AI’s data-driven computing paradigm because, under this paradigm, developers do not directly change the AI’s logic rules – i.e. Model. Instead, they indirectly affect the model by changing its training Dataset. This is what allows AI to scale and generalize far beyond what traditional compute could achieve: Instead of being encoded by developers, the training process simultaneously infers features in the training dataset and correlates them with their expected outputs – i.e. the data’s labels. The AI’s accuracy and ability to generalize is therefore a function of the quality and size of its training dataset, and of its training algorithm. This allows AI to scale and generalize at a level far beyond what traditional compute could achieve. However, because developers do not directly change the model, it also requires a significant shift in development methodologies.
Consider, for example, some of the techniques used by developers and adversaries to affect the model. Developers use deep analysis of the model and dataset to guide dataset changes or guardrail selection [2]. This deep analysis can include Explainable-AI (XAI) techniques, which extrapolate the model’s learned features and relationships [22]. Adversarial techniques – which are similar in spirit – can reveal key weights or activations to affect model behavior around specific data or features [8, 28, 17, 23].
This section presents a high level overview of the relationships of AI assets spanning across three representative stages:
@@ II Relationships between AI Assets
    Dataset Collection & Assembly (Figure 3)
@@
    Training (Figure 4)
@@
    Deploy & Inference (Figure 5)
@@
Each stage consists of multiple processes, depicted in the left-side column of each figure. Common operations or components within each process are denoted in the right-side column. Note that these stages are illustrative – it is not common for all operations to be implemented within a single monolithic system or process [24]. Circular transitions between processes or stages are also common. For example, within the Training stage, the Training and Model Analysis processes interleave many times before the AI is deemed ready to be deployed. Model Analysis could also reveal a need for additional data – requiring additional instances of Dataset Collection & Assembly. Accordingly, different variants of stage – or even of each process or operation – may exist within development and deployment pipelines. For example, developers add their own variants of the Dataset Collection & Assembly and Training stages when fine-tuning a pre-trained model.
Sections II-AII-C, and II-B describe each stage, and the relationships between various AI assets within them. Section II-D then describes various noteworthy security implications of these relationships.
II-A Dataset Collection & Assembly
Dataset Collection & Assembly encompasses the operations for collecting Raw Data and assembling it into a Dataset. Common operations are outlined in Figure 3.
Types of assets, and their relationships, that exist within this stage include:
  • Raw Data (Consumed, Modified). Raw Data can be in the analog or digital domains. It may also be processed before consumption, outside project scope. For example, images collected from a public page may have been resized upon upload. Raw Data has various intermediate states as it undergoes Pre-Processing and Transformation.
  • Dataset (Produced). The dataset is assembled from raw data which has been encoded, transformed, and labeled.
  • Validation Results (Produced, Consumed). Validation results influence subsequent iterations of Dataset Collection & Assembly. That is, they can cause changes to feature selection or augmentation, or motivate collection of additional data. They can also be useful to an adversary, as they reveal data types – and therefore features – the AI may have overfit on, may not recognize, or bias it may have internalized.
@@ II-A Dataset Collection & Assembly
II-B Training
Training encompasses the operations through which a Model learns from a training Dataset. Common operations are outlined in Figure 4.
Types of assets, and their relationships, that exist within this stage include:
  • Datasets (Consumed, Transiently Modified). Datasets are consumed from the Dataset Collection & Assembly stage. These include the Training Datasets from which the model will extract features and infer input/output relationships, and the Validation Datasets which are used to produce validation results. It is common practice for training data to be dynamically – and transiently – augmented [6].
  • Validation Results (Produced, Consumed). Validation results characterize the model’s performance and behavior, serving as guideposts for changes to training or other stages. They could, for example, motivate the use of specific guardrails (security mechanisms that constrain AI behavior within predefined boundaries) during Deploy & Inference. To an adversary, they can serve as a guide for attacks by revealing corner cases, unintended behaviors, or biases.
  • Model (Produced). Once the model’s behavior meets predetermined criteria, training is considered complete and the model is finalized; this criteria could be considered as a separate asset. It can subsequently be deployed for inference or used as a foundational model, for fine-tuning. Model checkpoints may also be produced and tracked, for provenance.

### v2

@@ I Introduction
This paper introduces an asset-centric methodology for threat modeling AI systems that addresses the unique security challenges posed by integrated AI agents. Unlike existing top-down frameworks that analyze individual attacks within specific product contexts, our bottom-up approach enables defenders to systematically identify how vulnerabilities—both conventional and AI-specific—impact critical AI assets across distributed infrastructures used to develop and deploy these agents. This methodology allows security teams to: (1) perform comprehensive analysis that communicates effectively across technical domains, (2) quantify security assumptions about third-party AI components without requiring visibility into their implementation, and (3) holistically identify AI-based vulnerabilities relevant to their specific product context. We have developed a threat analysis engine to automate holistic vulnerability identification (3). This approach is particularly relevant for securing agentic systems with complex autonomous capabilities. By focusing on assets rather than attacks, our approach scales with the rapidly evolving threat landscape while accommodating increasingly complex and distributed AI development pipelines. To facilitate security analysis, this work also identifies eight fundamental types of AI assets and maps out the range of ways in which adversaries can impact them.
@@ I-A AI-based Vulnerabilities
AI represents a shift in our computing paradigms towards data driven compute. Under traditional compute, a developer would create a program – i.e. logic rules – to generate a desired output from unseen input data. AI’s computing paradigm instead analyzes massive amounts of data to infer – i.e. learn or create – logic rules. These inferred rules are then used to generate a desired output when applied to unseen input data. The AI’s underlying decision making logic is, therefore, algorithmically inferred rather than explicitly encoded [9, 21]. This data driven nature is why AI’s decisions are often difficult to explain, and why AI is difficult to validate. Evasion, or jailbreak, attacks are an example of adversaries finding and exploiting corner cases or incorrectly correlated features within the AI’s inferred logic; i.e. the model. To compromise an AI’s decision, however, adversaries are not constrained to direct attacks on the model [28, 17, 22, 8]. Adversaries are also able to introduce vulnerabilities into an AI by affecting its training data, training process or algorithm, and validation process or goals. To augment their security policies against emerging AI-based vulnerabilities, defenders must reason about the many ways in which adversaries can affect their AI. This work enables such analysis by identifying eight fundamental asset types and mapping out the range of ways in which adversaries can impact them.
@@ I-C The case for a bottom-up perspective
Section II introduces common types of AI Assets and maps the range of potential capabilities an adversary may obtain over each. Section III illustrates how AI Assets are built and used throughout development and deployment pipelines. Section IV elaborates on the challenges defenders face when threat modeling these AI pipelines. Section V provides background on asset-centric threat modeling. Section VI describes our bottom-up, asset-centric, approach for threat modeling AI. Sections VII and VIII introduce Enterprise RAG and apply this analysis to it, as a case study. Section X describes existing frameworks for threat modeling AI and why they naturally complement an asset-centric approach. Section XI describes our ongoing work and next steps. Finally, Section XII concludes.
II Mapping AI Assets and Potential Capabilities
Defenders must adapt their security analysis to account for nuanced interactions between AI assets, stemming from AI’s data-driven computing paradigm. Under this paradigm, developers do not directly change the AI’s logic rules – i.e. the Model. Instead, they indirectly affect the model by changing the assets used to derive it. It is critical, therefore, that defenders understand the different types of AI assets and the ways in which adversaries might interact with them.
This section first identifies the types of AI assets (Section II-A), then expands conventional CIA notions into finer-grained capabilities (Section II-E), and finally presents various examples (Section II-F).
II-A Types of AI Assets
This section introduces eight fundamental types of AI assets. Each type represents a critical component that adversaries may target to compromise AI systems.
Inputs. Data or information to be processed by the AI. Inputs from benign users often require protection and isolation from other inputs. Most adversarial AI techniques assume that attackers can submit inputs to the AI – acting as malicious users.
Common Input components include: User Input: The primary data or query from the user Prompt Context: Historical conversation or session data Prompt Augmentations: Additional data like RAG outputs or system prompts Tool/API Descriptions: Metadata about available tools and their capabilities Tool/API Results: Outputs from external tool calls that augment the prompt
Outputs. Response(s) generated by the system. Increasingly, outputs are consumed by another model or agent. The output is, naturally, inherent to the functionality of the system.
Common Output components include: Generated Text/Content: The primary AI response Error Messages: System feedback that may leak information Tool or API Actions: Requests initiated by the AI
Output Details or Explanations. Additional details about the output, describing why the system provided or generated it. Some are generated as part of inference (confidences) and others require additional evaluation (detected features). Output details are distinguished from Outputs because adversarial techniques often require specific details or explanations.
Common Output Details or Explanations components include: Confidence Scores: Numerical certainty measures Feature Attributions: Which input features influenced the output Decision Paths: Step-by-step reasoning chains
Dataset. AI is data-driven compute, so the dataset is the most critical and sensitive asset of any AI/ML system. Common variants include: Training, Validation, Augmentation (RAG).
Datasets are composed of: Raw Data: Original, unprocessed information Encoded Data: Transformed representations (e.g., embeddings) Labels: Ground truth annotations for supervised learning Metadata: Provenance, timestamps, and quality indicators
Model Parameters: Weights (including biases), Activations, Updates. The static and dynamic data values which make up the model. Model theft, or extraction, seeks to approximate the model parameters.
Key Model Parameter components include: Weights: Learned parameters from training Activations: Intermediate computational states during inference Updates: Gradients or parameter changes during training/fine-tuning
Validation or Monitoring Criteria. Validation is a critical aspect of AI which increasingly relies on third party tools, services, and actors. Validation involves discovery of bias and corner cases in the model and training data. It’s common practice to validate the dataset during Data Collection & Assembly – before incurring training costs. Once deployed, model performance is also monitored; allowing developers to understand model performance on real-world data.
Validation or Monitoring Criteria examples include: User Input: The primary data or query from the user Prompt Context: Historical conversation or session data Prompt Augmentations: Additional data like RAG outputs or system prompts Tool/API Descriptions: Metadata about available tools and their capabilities Tool/API Results: Outputs from external tool calls that augment the prompt
Validation or Monitoring Results. Validation findings motivate additional data gathering or augmentation, changes to hyperparameters, or inform Guardrail selection. Validation results contain highly valuable knowledge for any bad actor, as bias and corner cases can serve as invaluable guides for subsequent attacks. Monitoring results also inform data collection and re-training.
Typical validation or monitoring results include: Performance Reports: Accuracy, precision, recall metrics Error Analysis: Patterns of model failures Drift Detection: Changes in model behavior over time
Hyperparameters: Model, Data, Training. Many hyperparameters exist throughout an AI System. They are related to the model design, data processing, or training configuration.
Hyperparameter categories include: Model Architecture: Layer types, sizes, connections Data Processing: Normalization methods, augmentation strategies Training Configuration: Learning rates, batch sizes, optimization algorithms
II-B Components
Each asset type is assembled from individual components that can be reasoned about separately for security analysis. For example, a dataset is made up of many individual pieces of data which have been encoded, transformed, and labeled. Common and example components are discussed throughout this work. During security analysis, components are assets themselves. However, the components of an asset type are ultimately a function of an AI’s functionality, implementation, and use within a specific system. To improve the generalizability and scalability of our vulnerability and attack analysis (Section VI), therefore, our approach relies on the affected asset types and not on their individual components. This allows defenders to reason about the components of each affected asset type based on their unique presentations within their system.
Key Principle: Capabilities over components translate to (typically weaker) capabilities over the whole asset. Gaining a capability over a component allows an adversary to gain a capability over the asset as a whole – however, the capability obtained over the whole asset is typically weaker than that gained over the individual component. Prompt injection attacks epitomize this concept. MCP Exploits, for example, are enabled by an adversary’s ability to Make Arbitrary Changes to MCP tool descriptions, which are used as an Input component. This provides the adversary with the ability to Contribute to the Input [26]. That is, the adversary lacks the ability to directly read or write to the Input as a whole, but is able to add data to it through tool descriptions. An adversary may ultimately need to compromise multiple components to obtain a desired capability over an asset type.
II-C Dependencies
Asset types exhibit dependencies that allow capabilities to propagate across the AI lifecycle. Obtaining a capability over an asset’s dependency allows an adversary to impact the downstream, dependent, asset – albeit indirectly and typically with a reduced capability.
Common dependency chains include:
  • Training Dataset → Model: The ability to make arbitrary changes to the training dataset allows an adversary to influence the trained model
  • Input → Output: The ability to make arbitrary changes to an inference input allows an adversary to influence the output
  • Validation Results → Hyperparameters: Manipulated validation results can influence hyperparameter selection
  • Model → Output Details: A compromised model affects all explanations and confidence scores
This dependency analysis enables modular security reasoning: defenders can focus on assets within their scope while understanding how upstream compromises might affect them. A defender’s security analysis can focus solely on the asset types that exist within a specific scope, or context, because downstream defenders can translate their analysis; quantifying the impact to dependent assets within downstream scopes or contexts. This greatly facilitates security analysis as dependencies or dependent assets may not even exist within a defender’s limited scope. Furthermore, identifying impacted dependencies allows defenders to reason about an attack’s root cause.
II-D Relationships
Asset types also exhibit relationships across models or agents in compound AI systems. The simplest of these is when an Output is directly used as an Input to another system; for example, when generated content is consumed by a Guardrail. To accurately quantify adversarial capabilities, defenders must be careful to identify transitions across models and agents.
Common cross-agent relationships:
  • Embedding Model → LLM: The ability to make arbitrary changes to a prompt before it’s processed by an embedding model provides the ability to influence the embedding model’s output. Because the embedding model’s output is the LLM’s input, the adversary can influence the LLM input.
  • LLM → Guardrail: LLM outputs directly become guardrail inputs, transferring capabilities
  • RAG → LLM: RAG outputs augment user prompts, providing contribute capability to the combined input
  • Agent → Agent: In multi-agent systems, one agent’s output becomes another’s context
Section III builds on these definitions to describe the different ways in which assets are built and used throughout development and deployment pipelines.
To enable security analysis, Section II-E expands conventional notions of Confidentiality, Integrity, and Availability (CIA) into a finer-grained range or potential capabilities.
II-E Potential Capabilities
To quantify the nuanced ways in which adversaries can manipulate AI Assets, this work expands the conventional notions of Confidentiality and Integrity.We identify nine potential capabilities. Both Confidentiality and Integrity are expanded into four capabilities. Availability is not expanded.
Confidentiality. We expand Confidentiality into four distinct capabilities which refer to an adversary’s ability to read an asset.
@@ II-E Potential Capabilities
    Inspect (Unlimited Read). This capability is akin to the conventional definition of Confidentiality, where an adversary gains the ability read the entire asset.
@@
    Partially Inspect (Limited Read). Refers to an adversary’s ability to read a subset of an asset.
@@
    Indirectly Inspect (Indirect or Inferred Read). Refers to an adversary’s ability to infer properties of an asset based on its usage or some other observable outcome. Adversaries may be able to reconstruct parts of the asset from this information – i.e. obtaining the ability to partially inspect it. Side channels often provide this capability.
  4. 4.
    Monitor. Refers to an adversary’s ability to observe whether an asset is consumed, and how often. The adversary, however, may lack the ability to directly or indirectly inspect the asset.
Integrity. We expand Integrity into four distinct capabilities which refer to an adversary’s ability to write to an asset.
  1. 1.
    Make Arbitrary Changes (Unlimited Write). This capability is akin to the conventional definition of Integrity, where an adversary gains the ability to modify the entire asset. That is, the adversary can change any part of the asset, without constraints. Adversary’s may self-limit to ensure that the asset retains functionality; e.g. remains the same data type.
  2. 2.
    Make Limited Changes (Limited Write). Refers either to an adversary’s ability to change a subset of an asset or to constraints on their changes; often, to both. Even with the ability to make arbitrary changes, adversaries must often ensure that any modifications are imperceptible or hidden [35, 30].
  3. 3.
    Influence (Indirect Write). Refers to an adversary’s ability to indirectly affect an asset. This capability is typically obtained by affecting a related asset. An adversary may also obtain this capability by exploiting fault injection vulnerabilities.
  4. 4.
    Contribute (Write without Read). This capability is uniquely prevalent, and devastating, for AI. It refers to an adversary’s ability to add data into an asset without being able to read the asset. This is the capability which enables prompt injections and some poisoning [31] or backdoor [30] attacks.
Availability. We do not expand the notion of availability beyond its conventional definition.
  1. 1.
    Availability (Withhold). Refers to an adversary’s ability to prevent access to an asset. Often achieved through a denial of service or by impacting integrity; e.g. deleting an asset. Throughout this work, we use withhold when referring to adversarial impact on availability.
As with conventional CIA, adversaries can also leverage one capability to obtain another. For example, consider a model fingerprinting attack [36]. First, an adversary leverages a timing channel to profile the model’s execution time; indirectly inspecting side effects of the model architecture. The adversary then correlates timing information to infer the model architecture; obtaining the ability to inspect it. Critically, the second step is necessary for the indirectly inspected information to become useful. Constant time programming techniques, for example, mitigate leakage through timing side channels by ensuring that no such correlation exists.
II-E1 Implied Capabilities, for Vulnerabilities
Figure 3: Obtaining certain capabilities (rows) provides adversaries with weaker, related, capabilities (columns). Capabilities are ordered from strongest (top-left) to weakest (bottom-right). Capabilities related to confidentiality are denoted in blues while those related to integrity are denoted in purple. A lack of capabilities – i.e. cannot inspect – is also included for clarity.
To facilitate analysis, Figure 3 also defines which capabilities are implicitly granted alongside the capability provided by a vulnerability. That is, if a vulnerability provides a strong capability, it also provides weaker ones. Notably, these definitions are designed to be conservative. This allows defenders to decide whether a specific capability is out of scope instead of analyzing each weaker capability to determine whether they are in scope.
The most straightforward examples are the ability to make limited changes by possessing the ability to make arbitrary changes or to partially inspect by possessing the ability to inspect. More subtle is an implied ability to influence or to contribute. It’s important to note that attackers can perform an attack which requires that they influence or contribute to an asset because they possess a stronger capability. As the attack surface providing the stronger capability is likely different, attackers will likely need to adapt their technique for that attack point and stronger capability.
Note that the notion of implied capabilities only exists for vulnerabilities. When describing how an adversary must compromise an asset, for an attack, the minimum capability is used. This is particularly relevant to the knowledge base discussed in Section VI-C. It does not make sense to reason about weaker capabilities when analyzing an attack because those capabilities would not enable the attack. Intuitively, this property is often seen in the offensive security literature: attack variants which require weaker capabilities – i.e. threat models or assumptions – are published as novel attacks. However, obtaining a stronger capability than required does enable the attack.
II-F Examples: Applying Capabilities to Real Scenarios
This section demonstrates how the capability model applies to real-world AI security scenarios, showing how capabilities over components translate to capabilities over complete assets.
II-F1 Example 1: Dataset Poisoning Through Multiple Sources
Consider a dataset $D$ composed of data from two sources: $D_{1}$ (adversary-controlled) and $D_{2}$ (trusted source).
Scenario: An adversary controls the data source for $D_{1}$ but their malicious modifications must pass manual inspection during labeling [30].
Capability Analysis:
  • Over $D_{1}$: The adversary has Make Limited Changes because modifications must remain imperceptible to pass inspection
  • Over $D_{2}$: The adversary has Cannot Inspect - no access to this trusted source
  • Over complete dataset $D$: The adversary has Contribute - they can add malicious data through $D_{1}$ without seeing $D_{2}$
Security Insight: This example demonstrates how capabilities over components translate to capabilities over the whole asset. Even without full dataset access, an adversary can still poison the training process.
II-F2 Example 2: Prompt Injection via Tool Descriptions
Consider a User Prompt input composed of multiple components in an LLM with tool access.
Scenario: An adversary exploits a vulnerability in how tool descriptions are processed [26].
Capability Analysis:
  • Over tool descriptions: Adversary has Make Arbitrary Changes
  • Over user’s original prompt: Adversary has Cannot Inspect
  • Over complete prompt (after augmentation): Adversary has Contribute
Attack Path: The adversary injects malicious instructions through tool descriptions, which get concatenated with the user’s legitimate prompt, effectively hijacking the LLM’s behavior without seeing the original query.
II-F3 Example 3: Model Extraction Through API Access
Consider an LLM deployment where users can query the model through an API.
Scenario: An adversary seeks to steal the model by observing input-output pairs.
Capability Analysis:
  • Over Model Parameters: Adversary has Indirectly Inspect through API responses
  • Over Inputs: Adversary has Make Arbitrary Changes (can craft any query)
  • Over Outputs: Adversary has Inspect (sees all responses to their queries)
  • Over Output Details: Depends on API configuration - may have access to confidence scores
Security Insight: The combination of arbitrary input control and output observation enables model extraction attacks, even without direct parameter access.
Key Takeaway: Capabilities degrade but propagate across model boundaries. Understanding these transitions is crucial for comprehensive security analysis.
III Processing and Using AI Assets
This section provides a comprehensive analysis of how assets and their associated capabilities evolve through the AI lifecycle. AI assets take different forms as they are built and used across AIOps stages – affecting the capabilities adversaries can obtain. This section identifies the unique set of assets, and potential capabilities, which exist across three representative AIOps stages:
  1. 1.
    Dataset Collection & Assembly. (Section III-A)
  2. 2.
    Training. (Section III-B)
  3. 3.
    Deploy & Inference. (Section III-C)
@@ III Processing and Using AI Assets
Key Insight: The same asset may offer different adversarial capabilities depending on the stage and process accessing it. Understanding these variations is crucial for comprehensive threat modeling. Common processes and operations are also identified for each illustrative stage. In practice, it is not common for all stages, processes, or even operations to be implemented as a single monolithic block [23]. Pipelines also often exhibit circular transitions between processes and stages. For example, within the Training stage, the Training and Model Analysis processes will interleave many times before the AI is deemed ready to be deployed. Model Analysis could also reveal a need for additional data – requiring an additional iteration through a Dataset Collection & Assembly stage. It’s also common to encounter unique variants of a stage, process, or operation. For example, developers must add their own variants of the Dataset Collection & Assembly and Training stages to fine-tuning a pre-trained model.
Due to space constraints, this section performs a deep dive into Dataset Collection & Assembly and provides a high level overview of both Training and Deploy & Inference.
III-A Dataset Collection & Assembly
@@
Dataset Collection & Assembly encompasses the processes for collecting Raw Data and assembling it into a Dataset. Processes and operations for this stage are summarized in Figure 4. We distinguish between processes (high-level activities like "Data Selection") and operations (specific implementations like "Filtering" or "Quality Control"). Assets present within this stage are:
  1. 1.
    Dataset: Composed of processed Raw Data components.
  2. 2.
    Hyperparameters: Configuration for data selection, pre-processing, transformation
  3. 3.
    Dataset Validation Criteria: Rules and thresholds for data quality
  4. 4.
    Dataset Validation Results: Outputs from validation processes
The Dataset is assembled from Raw Data components which have been processed; i.e. encoded, transformed, and labeled. As discussed in Section II-B, a capability over some raw data assets translates as a weaker capability over the dataset as a whole. For example, making arbitrary changes over a subset of raw data translates into making limited changes to the dataset. Common processes and operations for this stage are outlined in Figure 4.
Reasoning about Processes. Although the implementation of each process is unique to each pipeline, compromising the Collection, Data Selection, Pre-Processing, or Data Transformation processes typically allows an adversary to affect any property of raw data. This is because each of these processes is inherently designed to modify the data it consumes. Any impact to Confidentiality, Integrity, or Availability propagates when subsequent processes consume this modified data.
  • Collection consumes raw data from a source external to the system.
    Outputs raw data.
  • Data Selection consumes raw data and data selection hyperparameters. This processes discards raw data which does not meet selection criteria.
    Outputs selected data.
  • Pre-Processing consumes selected data and pre-processing hyperparameters.
    Outputs pre-processed data.
  • Transformation consumes pre-processed data and transformation hyperparameters.
    Outputs transformed data.
Certain processes inherently limit the capabilities which can be obtained. Dataset Validation, for example, loads the dataset but does not need to write it back. Exploits affecting dataset integrity within this process, therefore, would not propagate. However, an adversary could compromise the validation criteria or results to influence the dataset – exploiting a dependency chain. This same inherent limitation typically applies to all hyperparameters consumed within this stage. For capabilities to be permanent, instead of transient, adversaries must instead compromise these assets outside these processes.
  • Dataset Validation consumes an assembled dataset and validation criteria.
    Outputs validation results.
Affecting processes through Operations. To compromise a specific process, adversaries leverage the attack surfaces exposed by the operations implementing the process. Due to space constraints, we only provide examples for a small set of operations within Collection. The following are example attack surfaces exposed as raw data is collected through a camera sensor on an edge device. Collected data is buffered in local storage before being transferred to a centralized data storage server; e.g. an S3 bucket The other processes within Dataset Collection & Assembly are applied from this server. Attack surfaces are grouped based on the operation which exposes them. The specific vulnerabilities exploited at each surface would depend on that operation’s implementation and may exist at any point across its hardware/software stack.
  • Analog vs Digital Domain. An adversary might affect data within the analog domain, before it is collected by the camera sensor. As collected data will undergo selection and pre-processing, adversaries are constrained to making limited changes.
  • Collection Mechanism. An adversary might affect the availability of data by disabling the camera sensor. They might also make limited changes to the camera’s output by exploiting a vulnerability in the camera’s hardware implementation or drivers.
  • Retrieval & Transmission. An adversary might make limited changes to collected data while it’s buffered in local storage or intercept it in transit from the data collection endpoint to the central server.
III-B Training